OpenStudy (anonymous):

What is the taxidistance between (4, 7) and (17, 2)?

OpenStudy (merchandize):

The taxi-distance between (x₁ y₁) and (x₂, y₂) is |x₂ - x₁| + |y₂ - y₁|. Thus, the taxi-distance between (4,7) and (17,2) is: D = |17 - 4| + |2-7| = 13-5=8